Table 1            Estimated Total Returns                                                                                                 
State of the Economy Probability T-Bond SETX Golden S&P 500
Recession 5% 5% -19% 20% -14%
Below Average 15% 5% 2% 13% 3%
Average 45% 5% 9% 10% 11%
Above Average 25% 5% 34% 5% 22%
Boom 10% 5% 25% -5% 33%

Golden expects to pay $0.85 in dividends next year. Golden paid $0.67 in dividends 5 years ago. What should Golden sell for? If it sells for $14, should you buy? If Value Line says the growth rate should be 5.4%, what should Golden sell for? What is the required return using the constant growth model based on a $14 price and 5.4% growth rate?
Answer #1

Return Rate of Golden = 0.05(0.20) + 0.15(0.13) + 0.45(0.10) + 0.25(0.05) + 0.10(-0.05)

Return Rate of Golden = 8.20%

Growth Rate = (0.85/0.67)1/6 - 1 = 4.04%

Using DDM Model,

Stock Price of Golden = 0.85/(0.082 - 0.0404) = $20.43

One should buy golden at $14 as Intrinsic Value is $20.43